Lab Determination of a Rate Lavw Pre-Lab Questions 1. Determine the units of k for a first order reaction by solving the rate expression for k, and then substituting time units in seconds and concentrations in moles/liter Repeat this determination for a second order reaction 2. If the S,O,- was omitted, would the blue-black color appear in the reaction mixtures? What is the reason for adding S,0,- in the determination of the rate of the reaction between I- and S O,2-7 3. If2.0 mL of starch solution are added instead of 1.0 mL, what affect will this have upon the time for the blue-black color to appear? Explain. 4. If the reaction time for one of your runs in this experiment is 36 sec, calculate the reaction rate. Hint: 36 sec is the time required for a given change in concentration of S,0-
media%2F846%2F8461f0c0-6e4a-41a5-b927-04
media%2Ff34%2Ff344a7ec-35ce-4d0f-90b3-2c
Answer #1

For the first order reaction, the rate law expression is

Rate-kReactant

Rearranging the above equation

k=\frac{Rate}{\left [ reactant\right ]}

Therefore,

Unit of k=\frac{Unit of Rate}{\left [ Unit of reactant Concentration\right ]}

Unit of k=\frac{\left ( moles/liter \right ) \left /( second \right )}{\left [ moles/liter\right ]}

Unit of k=/\left ( second \right )

Unit of k=s^{-1}

For the second-order reaction, the rate law expression is

Rate= k\left [ Reactant \right ]^{2}

Rearranging the above equation

k=\frac{Rate}{\left [ reactant\right ]^{2}}

Therefore,

Unit of k=\frac{Unit of Rate}{\left [ Unit of reactant Concentration\right ]^{2}}

Unit of k=\frac{\left ( moles/liter \right ) \left /( second \right )}{\left [ moles/liter\right ]^{2}}

Unit of k=liter/mole/ second

Unit of k=liter\cdot mole^{-1}\cdot s^{-1}